# A predictive model for bubble–particle collisions in turbulence

### v1

@@ Abstract
The modelling of bubble–particle collisions is crucial to improving the efficiency of industrial processes such as froth flotation. Although such systems usually have turbulent flows and the bubbles are typically much larger than the particles, there currently exist no predictive models for this case which consistently include finite-size effects in the interaction with the bubbles as well as inertial effects for the particles simultaneously. As a first step, Jiang and Krug (2025) proposed a frozen turbulence approach which captures the collision rate between finite-size bubbles and inertial particles in homogeneous isotropic turbulence using the bubble slip velocity probability density function measured from simulations as an input. In this study, we further develop this approach into a model where the bubble–particle collision rate can be predicted a priori based on the bubble, particle, and turbulence properties. By comparing the predicted collision rate with simulations of bubbles with Stokes numbers of 2.8 and 6.3, and particles with Stokes numbers ranging from 0.01 to 2 in turbulence with a Taylor Reynolds number of 64, good agreement is found between model and simulations for Froude number $Fr\leq 0.25$. Beyond this range of bubble Stokes number, we propose a criterion for using our model and discuss the model’s validity. Evaluating our model at typical flotation parameters indicates that particle inertia effects are usually important. Generally, smaller bubbles, larger particles, and stronger turbulence increase the overall collision rate.
@@
Fully predictive model of collision rate in homogeneous isotropic turbulence
Includes effects of particle inertia and bubble flow field distortion
Bubble, particle, and turbulence properties are the only inputs required
Potential subgrid scale model for industrial scale froth flotation simulations
Direct numerical simulations of finite-size bubbles and point-particles in turbulence
@@ 1 Introduction
Bubble–particle collisions lie at the heart of froth flotation — an industrial process which selectively recovers particles based on their hydrophobicity with applications ranging from mineral extraction to wastewater treatment (Rubio et al., 2002; Nguyen and Schulze, 2004). In this process, bubbles and particles are injected into a cell filled with water so they can collide with each other. As the bubbles and particles collide, the target particles (which are hydrophobic) attach to the bubbles and are floated to the surface, where they can be removed. In other words, every single particle has to collide with a bubble before it is recovered. Given the broad range of application of froth flotation, as well as the large size of flotation cells and the complex flows therein which precludes particle-resolved simulations, there is a strong incentive to model the bubble–particle collision rate. This parameter is usually reported in terms of the collision kernel
$\Gamma=Z/n_{b}n_{p},$ (1)
where $Z$ is the collision rate per unit volume and $n_{b}$ and $n_{p}$ are the bubble and particle number densities, respectively, to facilitate comparison between systems with a different number of particles.
1.1 Bubble–particle collisions in still fluid
A key characteristic of bubble–particle collisions in flotation of mineral particles lies in the huge size difference between bubbles and particles, with typical bubble-to-particle radius ratios $r_{b}/r_{p}\sim\mathrm{\textit{O}}(10)$ (Rubio et al., 2002; Ngo-Cong et al., 2018; Kostoglou et al., 2020a). This means that the actual collision rate is significantly affected by the distortion of the flow field due to the bubbles. Many related studies have focused exclusively on the case of a bubble rising in a still fluid with a constant slip velocity of $w_{b}$ (see e.g. Huang et al., 2012; Nguyen and Schulze, 2004; Sarrot et al., 2005; Schulze, 1989; Weber and Paddock, 1983; Zheng et al., 2024). In this limit, the problem is deterministic. As figure 1(a) illustrates, only particles that are initially (i.e. in the undisturbed flow far from the bubble) located within a distance $r_{g}$ from the centreline of the cylinder swept by the bubble eventually collide with the bubble ($r_{g}$ is hence known as the ‘grazing radius’). The ratio of the number of particles that actually hit the bubble to the number of particles inside the cylinder swept by the bubble is the ‘collision efficiency’
$E_{c}=\bigg{(}\frac{r_{g}}{r_{b}}\bigg{)}^{2},$ (2)
such that when the particle number density immediately upstream of the bubble is equal to the global value, the collision kernel is given by
$\Gamma=\pi r_{b}^{2}w_{b}E_{c}.$ (3)
@@ 1.1 Bubble–particle collisions in still fluid
In general, $E_{c}$ is determined by five nondimensional parameters, i.e.
$E_{c}=f\bigg{(}Re_{b}=\frac{2r_{b}w_{b}}{\nu},\frac{r_{b}}{r_{p}},St_{p}^{\prime}=\frac{\tau_{p}w_{b}}{2r_{b}},\frac{v_{s}}{w_{b}},\theta_{cap}\bigg{)},$ (4)
@@
where $Re_{b}$ is the bubble Reynolds number, $\nu$ is the kinematic viscosity, $St_{p}^{\prime}$ is the particle Stokes number based on the bubble slip velocity, $\tau_{p}=r_{p}^{2}(2\rho_{p}/\rho_{f}+1)/(9\nu)$ is the particle response time, $\rho_{p}$ is the particle density, $\rho_{f}$ is the fluid density, $v_{s}$ is the particle settling speed, and $\theta_{cap}$ is the bubble contamination angle. In the discussion below, we restrict ourselves to the case of fully contaminated bubbles ($\theta_{cap}=\pi$) and non-settling particles ($v_{s}=0$). The former is reasonable for bubbles in flotation (Nguyen and Schulze, 2004; Huang et al., 2012), while the latter assumption is strictly valid when $v_{s}$ is much smaller than the bubble velocity and is imposed here as a simplification of the real situation.
@@
In the $St_{p}^{\prime}\to 0$ limit particles perfectly follow the fluid flow, a collision occurs if the streamline carrying the particle comes within the collision distance $r_{c}$. This is termed the ‘interception mechanism’ (Schulze, 1989) and Sarrot et al. (2005) showed that
@@
where $a$ and $b$ are fitting parameters (see A). Schulze (1989) proposed to model the overall collision efficiency as a superposition of the interception and inertial contributions, leading to
$E_{c}=E_{c}^{(i)}+E_{c}^{(in)}\bigg{(}1-\frac{E_{c}^{(i)}}{(1+r_{p}/r_{b})^{2}}\bigg{)}.$ (7)
Here, the term in brackets ensures that $E_{c}$ scales as $(1+r_{p}/r_{b})^{2}$ in the high $St_{p}^{\prime}$ limit.
Figure 1: (a) Sketch of a particle with $St_{p}^{\prime}=0$ travelling along the grazing trajectory (black solid line) to collide with a bubble in still fluid. The red dashed line is the corresponding streamline. The collision cylinder used to define the collision efficiency is also shown. (b) Sketch of a bubble–particle collision with finite $St_{p}^{\prime}$ where the grazing trajectory deviates from the corresponding streamline.
1.2 Bubble–particle collisions in turbulence
The above deterministic picture, however, falls short in the presence of turbulence, which is an important factor in most applications. In these, turbulence is either induced by the rising bubbles or even generated in order to increase the bubble–particle collision rate (Jiang and Krug, 2025). Turbulence introduces characteristic time and length scales into the problem leading to the emergence of additional dimensionless numbers. The most relevant ones are the Taylor Reynolds number $Re_{\lambda}$, which indicates the turbulence intensity, the Froude number $Fr$ characterising the strength of turbulence relative to gravity, and the bubble and particle Stokes numbers based on the Kolmogorov time scale $St_{b}$ and $St_{p}$. These are defined as follows:
where $u^{\prime}$ is the single-component root-mean-square (r.m.s.) fluid velocity, $\nu$ is the kinematic viscosity, $\varepsilon$ is the average rate of turbulent dissipation, $\mathfrak{g}$ is the gravitational acceleration, $\tau_{b}=r_{b}^{2}(2\rho_{b}/\rho_{f}+1)/(9\nu)$ is the bubble response time (the subscript ‘b’ denotes bubbles), and $\tau_{\eta}=\sqrt{\nu/\varepsilon}$ and $u_{\eta}=(\nu\varepsilon)^{0.25}$ are the Kolmogorov time scale and velocity scale, respectively. Accounting for the effect of turbulence in addition to the flow distortion by the bubble is a long standing problem in the literature (Pyke et al., 2003; Nguyen et al., 2016; Wang et al., 2018; Hassanzadeh et al., 2018; Wang et al., 2020, 2022). One approach is to consider an extra subprocess for bubble–particle collisions: ‘geometric collisions’, which are encounters resulting purely from turbulence and buoyancy-driven bubble and particle motions neglecting any distortion to the flow field caused by the bubbles (Bloom and Heindel, 2002, 2003; Pyke et al., 2003; Koh and Schwarz, 2006). The total collision rate is then given by multiplying the geometric collision rate in turbulence with the collision efficiency $E_{c}$ in still fluid based on an ensemble-average value of $Re_{b}$. However, this is problematic as the instantaneous and averaged values of $Re_{b}$ can differ significantly. Additionally, Pyke et al. (2003) incorrectly defines $Re_{b}$ with the bubble rise velocity instead of the bubble slip velocity. In contrast, the recent model by Kostoglou et al. (2020a) implicitly includes the collision efficiency concept in turbulence in a consistent manner. This is done by directly evaluating the particle influx around an isolated bubble that is exposed to an unsteady incident flow whose fluctuations are determined by the bubble r.m.s. slip velocity, and by incorporating the distorted flow field around the bubble using the expression given by Nguyen and Schulze (2004). In addition, they consider the collisions caused by small scale turbulence. Their model is based on the combined surface and velocity integral
$\Gamma^{(Kos)}$ $=$ $\oiint_{S_{c}}\iiint\limits_{-\infty}^{\infty}\int_{0}^{\infty}\Lambda(-\Delta v_{r}\cdot f(w_{bo,x})f(w_{bo,y})f(w_{bo,z})f(u_{pt})$
$\mathrm{d}u_{pt}\mathrm{d}w_{bo,x}\mathrm{d}w_{bo,y}\mathrm{d}w_{bo,z})\mathrm{d}S,$
$\Lambda(x)=\left\{\begin{aligned} x\quad&\text{for $x>0$,}\\ 0\quad&\text{otherwise,}\\ \end{aligned}\right.$ (10)
$f(\cdot)$ is the probability density function (p.d.f.), $w_{bo,x,y,z}$ are components of the bubble slip velocity with respect to the fluid velocity in the absence of the bubble, and $u_{pt}$ is the relative bubble–particle velocity due to local velocity gradients (see B for the complete expression). The model is proposed for the case of uniformly distributed tracer-like particles colliding with bubbles so it has limited applicability in real-life, where most particles are inertial. Jiang and Krug (2025) therefore developed an approach that applies to inertial one-way coupled particles. To obtain the collision kernel, they decompose the bubble trajectory into tiny fragments, for which the bubble velocity can be considered constant. During the time it takes for the bubble to travel along such a fragment, they employ a ‘frozen turbulence’ approximation (i.e. the flow field close to the bubble is considered to be steady and uniform), such that the instantaneous process is equivalent to the deterministic still fluid case (described in §1.1) with a bubble slip velocity corresponding to the instantaneous value. The instantaneous collision efficiency and collision kernel are hence given by (7) and (3), respectively. Integrating this over the entire bubble trajectory results in the overall collision kernel
$\Gamma=\pi r_{b}^{2}\int_{0}^{\infty}E_{c}(Re_{b},St_{p}^{\prime})w_{b}f(w_{b})\mathrm{d}w_{b}.$ (11)
As $E_{c}(Re_{b},St_{p}^{\prime})$ is defined by (7) and can be parametrised, Jiang and Krug (2025) used the measured $f(w_{b})$ from their simulations to validate the underlying ‘frozen turbulence’ approximation for $St_{b}=11$. The missing ingredient to turn this approach into a predictive model is therefore a suitable model that predicts the p.d.f. of the bubble slip speed $f(w_{b})$ using flow and bubble parameters as inputs. To develop this is the aim of this article. We present a bubble–particle collision model which consistently accounts for turbulence and the distorted flow field around the bubble based on (11), and show that it is applicable for a wide range of $St_{p}$ and multiple values of $St_{b}$.
This paper is organised as follows: we present our modelling approach of $f(w_{b})$ and review the related models in §2, then we describe the simulations conducted in §3. The simulation results are compared with the model predictions in §4. This is followed by a discussion on our model’s regime of validity and its practical implications in §5. Finally, the conclusions are drawn in §6.
2 Modelling the bubble slip velocity probability density function
To model the probability density function of $w_{b}$, we consider normally distributed bubble slip velocity components $w_{b,i}$ ($i=x,y,z$), i.e.

### v2

@@ Abstract
The modelling of bubble–particle collisions is crucial to improving the efficiency of industrial processes such as froth flotation. Although such systems usually have turbulent flows and the bubbles are typically much larger than the particles, there currently exist no predictive models for this case which consistently include finite-size effects in the interaction with the bubbles as well as inertial effects for the particles simultaneously. As a first step, Jiang and Krug (2025) proposed a frozen turbulence approach which captures the collision rate between finite-size bubbles and inertial particles in homogeneous isotropic turbulence using the bubble slip velocity probability density function measured from simulations as an input. In this study, we further develop this approach into a model where the bubble–particle collision rate can be predicted a priori based on the bubble, particle, and turbulence properties. By comparing the predicted collision rate with simulations of bubbles with Stokes numbers of 2.8 and 6.3, and particles with Stokes numbers ranging from 0.01 to 2 in turbulence with a Taylor Reynolds number of 64, good agreement is found between model and simulations for Froude number $Fr\leq 0.25$. Beyond this range of bubble Stokes number, we propose a criterion for using our model and discuss the model’s validity. Evaluating our model at typical flotation parameters indicates that particle inertia and settling effects are usually important. Generally, smaller bubbles, larger particles, and stronger turbulence increase the overall collision rate.
@@
Developed a fully predictive model for the bubble-particle collision rate in homogeneous isotropic turbulence
Incorporates particle inertia and bubble-induced flow field distortion into the collision model
Requires only bubble, particle, and turbulence properties as model inputs
Provides a potential subgrid-scale model to support industrial-scale froth flotation simulations
Validated through direct numerical simulations of finite-size bubbles and point-particles in turbulent flows
@@ 1 Introduction
Bubble–particle collisions lie at the heart of froth flotation — an industrial process which selectively recovers particles based on their hydrophobicity with applications ranging from mineral extraction to wastewater treatment (Rubio et al., 2002; Nguyen and Schulze, 2004). In this process, bubbles and particles are injected into a cell filled with water so they can collide with each other. As the bubbles and particles collide, the target particles (which are hydrophobic) attach to the bubbles and are floated to the surface, where they can be removed. In other words, every single particle has to collide with a bubble before it is recovered. Given the broad range of application of froth flotation, as well as the large size of flotation cells and the complex flows therein which precludes particle-resolved simulations, there is a strong incentive to model the bubble–particle collision rate. The collision rate per volume $Z$ can be expressed as:
$Z=\Gamma n_{b}n_{p},$ (1)
where $n_{p}$ and $n_{b}$ are the particle and bubble number density. The collision kernel $\Gamma$ serves as a fundamental parameter in quantifying the frequency at which bubbles and particles collide per unit volume and time.
In quiescent fluid, the bubble–particle collision process is primarily deterministic, governed by hydrodynamic mechanisms such as interception, gravitational sedimentation, and inertia. Particles approach rising bubbles along trajectories shaped by these mechanisms, and the collision kernel in such systems can be equivalently quantified by the collision efficiency ($E_{c}$). The collision efficiency quantifies the probability of collision, defined as the ratio of the number of particles that actually collide with the bubble to the total number of particles encountered along the bubble’s trajectory. The dominant collision mechanism depends strongly on particle properties, notably density $\rho_{p}$ and radius $r_{p}$, relative to those of the bubble. Typically, particles are significantly smaller than bubbles, making the local flow field around the bubble crucial for the collision dynamics. In the low-inertia limit, particles closely follow fluid streamlines, resulting in interception-driven collisions. The interceptional collision efficiency can be analytically derived based on the flow solutions around a bubble, as presented by Weber and Paddock (1983); Yoon and Luttrell (1989); Sutherland (1948), and describes how streamlines bring particles into contact with the bubble surface. When particle inertia increases, particles deviate from streamlines and collide due to inertial effects as well, leading to a marked rise in collision efficiency (Schulze (1989)). Additionally, particle gravitational settling can enhance collision rates, especially for dense or large particles, with corresponding models extending the efficiency predictions (Ralston et al. (1999); Weber and Paddock (1983)). Intuitively, the contributions of interception, inertia, and gravity to the total collision efficiency may be considered additive (Schulze (1989)). More recently, Huang et al. (2012) studied the impact of the interface contamination on the collision efficiency by means of direct numerical simulations.
While the deterministic mechanisms governing bubble–particle collisions in quiescent fluid provide valuable foundational insights, real flotation systems typically operate under turbulent conditions where flow unsteadiness and velocity fluctuations profoundly influence collision dynamics. Unlike the relatively predictable trajectories in static environments, turbulence introduces stochastic relative motions and enhanced mixing that can significantly increase collision rates beyond those captured by quiescent fluid models. To this end, more parameters are relevant in determining the collision kernel in turbulence. The most relevant ones are the Taylor Reynolds number $Re_{\lambda}$, which indicates the turbulence intensity, the Froude number $Fr$ characterising the strength of turbulence relative to gravity, and the bubble and particle Stokes numbers based on the Kolmogorov time scale $St_{b}$ and $St_{p}$. These are defined as follows:
where $u^{\prime}$ is the single-component root-mean-square (r.m.s.) fluid velocity, $\nu$ is the kinematic viscosity, $\varepsilon$ is the average rate of turbulent dissipation, $\mathfrak{g}$ is the gravitational acceleration, $\tau_{b}=r_{b}^{2}(2\rho_{b}/\rho_{f}+1)/(9\nu)$ is the bubble response time, $\tau_{p}=r_{p}^{2}(2\rho_{p}/\rho_{f}+1)/(9\nu)$ is the particle response time, $r_{b}$ is the bubble radius, $\rho_{b}$ is the bubble density, $\rho_{f}$ is the fluid density, and $\tau_{\eta}=\sqrt{\nu/\varepsilon}$ and $u_{\eta}=(\nu\varepsilon)^{0.25}$ are the Kolmogorov time scale and velocity scale, respectively. The particle response time characterizes the time scale over which a particle adjusts its velocity to match changes in the surrounding fluid velocity. It can be derived by nondimensionalizing the equation of motion for point particles subjected to fluid forces, as detailed in seminal works (Maxey and Riley, 1983; Gatignol, 1983). In turbulent flows, this formulation is generally valid for particles that are small compared to the Kolmogorov length scale $\eta$.
Modeling the bubble–particle collision kernel fundamentally depends on a detailed understanding of both bubble and particle dynamics within turbulent flows. Early models for bubble–particle collisions have naturally drawn upon the extensive literature on particle–particle collisions (see for example Pumir and Wilkinson (2016)), where dynamics are often described by point-particle approximations. One of the seminal works in this area is the model proposed by Saffman and Turner (1956), which accounts for shear-induced collisions between tracer-like particles in turbulence. In contrast, Abrahamson (1975) developed a collision kernel model applicable to particles with very large inertia, characterizing collision rates driven primarily by relative inertial motion. Practically, flotation systems operate with bubbles and particles that lie in an intermediate inertia regime. Addressing this, Yuu (1984) proposed a collision model that incorporates both shear and inertia induced relative motions for particles with intermediate Stokes numbers, thus bridging the gap between the two limiting cases. Nevertheless, these models predominantly quantify the “geometric collision” rate typical of particle–particle interactions (Chan et al. (2023) and Chan et al. (2025) for an overview). Insights from studies on bubble–particle collision efficiency in quiescent fluids clearly demonstrate that disregarding the pronounced size difference between bubbles and particles introduces significant inaccuracies in collision kernel estimates. The large size disparity indicates that the flow disturbance created by the bubble strongly influences particle trajectories, fundamentally altering collision probabilities compared to particle–particle interactions. However, accounting for the effect of turbulence in addition to the flow distortion by the bubble is a long standing problem in the literature (Pyke et al., 2003; Nguyen et al., 2016; Wang et al., 2018; Hassanzadeh et al., 2018; Wang et al., 2020, 2022). One approach has been to estimate the collision kernel in turbulent flows by multiplying the geometric collision kernel by the collision efficiency (Pyke et al. (2003); Koh and Schwarz (2006)). However, this approach is conceptually problematic as the instantaneous and averaged values of $Re_{b}$ can differ significantly. Additionally, Pyke et al. (2003) incorrectly defines $Re_{b}$ with the bubble rise velocity instead of the bubble slip velocity.
In contrast, the recent model by Kostoglou et al. (2020a) takes a different approach and implicitly includes the collision efficiency concept in turbulence in a consistent manner. This is done by directly evaluating the particle influx around an isolated bubble that is exposed to an unsteady incident flow whose fluctuations are determined by the bubble r.m.s. slip velocity, and by incorporating the distorted flow field around the bubble using the expression given by Nguyen and Schulze (2004). In addition, they consider the collisions caused by small scale turbulence. Their model is proposed for the case of uniformly distributed tracer-like particles colliding with bubbles, so it has limited applicability in real-life, where inertial effects are significant for most particles. More recently, Jiang and Krug (2025) developed an approach applicable to inertial, one-way coupled particles by decomposing the bubble trajectory into small segments over which the bubble slip velocity is assumed constant. This reduces the instantaneous collision process to a deterministic quiescent fluid scenario with a slip velocity equal to the instantaneous value. However, in their study the bubble slip velocity distribution is not predicted, but it is obtained directly from numerical simulations. As a result, the model remains unclosed in terms of collision kernel prediction. The present work builds upon the framework of Jiang and Krug (2025) by developing a complete model for the slip velocity probability distribution, thereby enabling standalone prediction of the bubble–particle collision kernel without reliance on additional simulation data.
This paper is organised as follows: We first We first present our model in §2, then we describe the simulations conducted in §3. The simulation results are compared with the model predictions in §4. This is followed by a discussion on our model’s regime of validity and its practical implications in §5. Finally, the conclusions are drawn in §6.
Figure 1: (a) Sketch of the ’frozen turbulence’ collision model along the bubble trajectory where it collides with particles at multiple instances. The dashed line represents streamlines around the bubble and the red solid lines denote particle trajectories under the frozen turbulence approximation. The sizes of the bubble, the particles and the particle trajectories are exaggerated for visibility.(b) Sketch of a particle with $St_{p}^{\prime}=0$ travelling along the grazing trajectory (black solid line) to collide with a bubble in still fluid. The red dashed line is the corresponding streamline. The collision cylinder used to define the collision efficiency is also shown. (c) Sketch of a bubble–particle collision with finite $St_{p}^{\prime}$ where the grazing trajectory deviates from the corresponding streamline.
2 Modelling approach
2.1 General concept
We start with a brief description of the collision framework proposed in Jiang and Krug (2025). To obtain the collision kernel, the bubble trajectory is decomposed into tiny fragments, for which the bubble velocity can be considered constant. During the time it takes for the bubble to travel along such a fragment, a ‘frozen turbulence’ approximation is employed (i.e. the flow field close to the bubble is considered to be steady and uniform), such that the instantaneous process is equivalent to the deterministic still fluid case with a bubble slip velocity corresponding to the instantaneous value. This approximation therefore focuses on the turbulence effects on bubbles instead of particles since for finite-size bubbles, the flow near the bubble deviates from homogeneous isotropic turbulence such that turbulence effects on particles may be less significant. Figure 1(a) sketches the frozen turbulence approximation for an example bubble trajectory, where the streamlines and particle trajectories around the bubble positions at three instants are shown. At every instant, the collision kernel can be evaluated by the collision efficiency with the corresponding conditions (see Jiang and Krug (2025) for details). Integrating this over the entire bubble trajectory results in the overall collision kernel
$\Gamma=\pi r_{b}^{2}\int_{0}^{\infty}E_{c}(Re_{b},St_{p}^{\prime})w_{b}f(w_{b})\mathrm{d}w_{b},$ (3)
where $f(\cdot)$ is the probability density function (p.d.f.) and $E_{c}$ represents the collision efficiency in a still fluid as a function of the bubble Reynolds number ($Re_{b}=2r_{b}w_{b}/\nu$) and the particle Stokes number ($St_{p}^{\prime}=\tau_{p}w_{b}/2r_{b}$), both defined with respect to the bubble slip velocity $w_{b}$. As illustrated in figure 1(b), only particles that are initially (i.e. in the undisturbed flow far from the bubble) located within a distance $r_{g}$ from the centreline of the cylinder swept by the bubble eventually collide with the bubble ($r_{g}$ is hence known as the ‘grazing radius’). The collision efficiency $E_{c}$ can therefore be expressed as the ratio between the area of a circle with radius $r_{g}$ to that of a circle with radius $r_{b}$ (Schulze, 1989; Dai et al., 1998):
[…]
In order to turn Eq.(3) into a predictive model, parametrisations for both $f(w_{b})$ and $E_{c}$ must be established. These will be presented in the following two sections, respectively.
2.2 Parametrisation of the collision efficiency
@@
In general, $E_{c}$ is determined by five nondimensional parameters (Sarrot et al., 2005; Huang et al., 2012; Schuch and Löffler, 1978; Schulze, 1989):
$Re_{b},\,\frac{r_{b}}{r_{p}},\,St_{p}^{\prime},\,\frac{v_{s}}{w_{b}},\,\theta_{cap},$ (5)
@@
which are the bubble Reynolds number, the bubble-particle size ratio, the particle Stokes number based on bubble slip velocity, the particle settling speed ratio, and the bubble contamination angle, respectively. Additionally, $v_{s}$ is the particle settling speed. In the discussion below, we restrict ourselves to the case of fully contaminated bubbles ($\theta_{cap}=\pi$), which is reasonable for bubbles in flotation (Nguyen and Schulze, 2004; Huang et al., 2012).
@@
In the $St_{p}^{\prime}\to 0$ limit particles perfectly follow the fluid flow. A collision then occurs if the streamline carrying the particle comes within the collision distance $r_{c}$. This is termed the ‘interception mechanism’ (Schulze, 1989) and Sarrot et al. (2005) showed that
@@
where $a$ and $b$ are fitting parameters, which were determined through additional simulations (see A). Furthermore, gravitational settling of heavy particles also affects the collision efficiency. To capture this effect, Weber and Paddock (1983) incorporated the particle settling velocity additively into the local fluid velocity at the collision position and integrated the resulting particle flux over the collision angle up to a critical value $\theta_{c}$, thereby deriving the gravitational collision efficiency:
$E_{c}^{(g)}=-\left(1+\frac{r_{p}}{r_{b}}\right)^{2}\frac{v_{s}}{w_{b}}\sin^{2}\theta_{c},$ (8)
where we adopt the expression of $\theta_{c}$ from Kostoglou et al. (2020a) (also provided in B). Note that since particle settling corresponds to negative values of $v_{s}$, the resulting gravitational contribution to the collision kernel is positive in this case. Schulze (1989) proposed to model the overall collision efficiency as a superposition of the interception, gravitational, and inertial contributions, leading to
$E_{c}=E_{c}^{(i)}{+E_{c}^{(g)}}+E_{c}^{(in)}\bigg(1-\frac{E_{c}^{(i)}}{(1+r_{p}/r_{b})^{2}}\bigg).$ (9)
Here, the term in brackets ensures that $E_{c}$ scales as $(1+r_{p}/r_{b})^{2}$ in the high $St_{p}^{\prime}$ limit. We note that the inertial and gravitational contributions are treated as additive, though this assumption is not strictly valid in every case. It is applied in this expression with the assumption that either $E_{c}^{(i)}+E_{c}^{(g)}$ or $E^{(in)}$ is dominant (Schulze, 1989). Nevertheless, this formulation represents, to the best of our knowledge, the most physically consistent and predictive expression currently available for describing the combined influence of particle inertia and gravity. Note that the particle gravitational settling effect is not included in the work of Jiang and Krug (2025). As the particle settling may have a role when the particle Stokes number is large, we here extend the model of Jiang and Krug (2025) by including $E_{c}^{(g)}$.
2.3 Modelling the bubble slip velocity probability density function
The final missing ingredient to turn the approach outlined in Eq.(3) into a predictive model is a suitable model that predicts the p.d.f. of the bubble slip speed $f(w_{b})$ using flow and bubble parameters as inputs. For this, we start with considering the slip velocity vector $\mathbf{w}_{b}$ with normally distributed velocity components $w_{b,i}$ ($i=x,y,z$ indicates the direction), i.e.